# appendixReferences

### v1

@@ Abstract
Endowing machines with abstract reasoning ability has been a long-term research topic in artificial intelligence. Raven’s Progressive Matrix (RPM) is widely used to probe abstract visual reasoning in machine intelligence, where models need to understand the underlying rules and select the missing bottom-right images out of candidate sets to complete image matrices. The participators can display powerful reasoning ability by inferring the underlying attribute-changing rules and imagining the missing images at arbitrary positions. However, existing solvers can hardly manifest such an ability in realistic RPM problems. In this paper, we propose a conditional generative model to solve answer generation problems through Rule AbstractIon and SElection (RAISE) in the latent space. RAISE encodes image attributes as latent concepts and decomposes underlying rules into atomic rules by means of concepts, which are abstracted as global learnable parameters. When generating the answer, RAISE selects proper atomic rules out of the global knowledge set for each concept and composes them into the integrated rule of an RPM. In most configurations, RAISE outperforms the compared generative solvers in tasks of generating bottom-right and arbitrary-position answers. We test RAISE in the odd-one-out task and two held-out configurations to demonstrate how learning decoupled latent concepts and atomic rules helps find the image breaking the underlying rules and handle RPMs with unseen combinations of rules and attributes.
@@ 1 Introduction
The abstract reasoning ability is pivotal to abstracting the underlying rules from observations and quickly adapting to novel situations Cattell (1963); Zhuo & Kankanhalli (2021); Małkiński & Mańdziuk (2022a), which is the foundation of various cognitive processes Gray & Thompson (2004) such as number sense Dehaene (2011), spatial reasoning Byrne & Johnson-Laird (1989), and physical reasoning McCloskey (1983). In machine intelligence, human-like abstract reasoning can help models apply acquired skills in unseen tasks Barrett et al. (2018). For example, generalize the understanding of collisions from triangles to squares. Therefore, endowing machine learning models with human-like abstract reasoning is the path to achieving higher-intelligence systems and the long-lasting research topic of artificial intelligence Chollet (2019); Małkiński & Mańdziuk (2022b).
@@
Raven’s Progressive Matrix (RPM) is a classical test to probe the abstract reasoning ability in human and machine intelligence Małkiński & Mańdziuk (2022a). In RPM tests, models need to choose one image out of eight candidates to fill the bottom-right corner of a 3$\times$3 image matrix Raven & Court (1998). Some studies suggest that participators can display powerful reasoning ability by directly imagining the missing images Hua & Kunda (2020); Pekar et al. (2020). In machine learning models, answer-generation tasks can more accurately reflect the understanding of underlying rules than discriminative ones Mitchell (2021). For example, some models can find shortcuts to discriminative tasks, that is, they select answers via candidate sets instead of the given context of RPMs.
@@
Many RPM solvers choose images by filling candidates and scoring the entire matrix and thus cannot imagine answers from the given context Barrett et al. (2018); Wu et al. (2020); Hu et al. (2021). Some generative solvers have been proposed to solve answer-generation tasks Pekar et al. (2020); Zhang et al. (2021b; a). They directly generate the bottom-right images and select answers by comparing the outcomes and candidates, which is a significant advance for generative abstract reasoning. However, some of them do not learn interpretable attributes and attribute-changing rules from the image Pekar et al. (2020). And the others usually introduce non-negligible prior information when representing attributes Zhang et al. (2021b; a). On the other hand, the training of these models still relies on the distractors in candidate sets, bringing the potential risk of being affected by the distribution of candidate sets Hu et al. (2021); Benny et al. (2021).
@@
Deep latent variable models (DLVMs) Kingma & Welling (2013) use different latent spaces to capture underlying structures of noisy observations Edwards & Storkey (2017); Eslami et al. (2018); Garnelo et al. (2018); Kim et al. (2019). Previous works have proved that generative RPM problems can be solved by regarding attributes and attribute-changing rules as latent concepts Shi et al. (2021), which can produce multiple solutions to a generative RPM problem by executing the generative process many times. These generative solvers take conditional generation processes for answer generation Sohn et al. (2015), and thus the distractors in candidate sets are not necessary for training. Although they illustrate the possibility of solving RPM tests with DLVMs, understanding complex discrete rules in realistic RPM problems and abstracting global rules from attribute-specific rules is still a challenge for DLVMs.
@@
This paper proposes a conditional generative model to solve generative RPM problems through Rule AbstractIon and SElection (RAISE). RAISE encodes the shared attributes of images (e.g., object size and shape) as a group of independent latent concepts to bridge the high-dimensional images and the knowledge of rules. The rules of RPMs are decomposed into atomic rules in terms of latent concepts, which are abstracted as a set of learnable parameters shared among RPMs. When predicting target images, RAISE picks up proper rules for each latent concept and composes them as the integrated rule of an RPM. The conditional generative process defined in RAISE indicates how to use the global knowledge of rules to imagine (generate) target images interpretably. The latent concepts are automatically learned by RAISE without meta information of image attributes to reduce the artificial priors in the learning process. RAISE can be trained under semi-supervised settings, requiring only a small amount of rule annotations to outperform the compared models in non-grid configurations. RAISE also supports the generation of missing images at arbitrary and even multiple positions. By predicting the target images at random positions of RPMs, the training process of RAISE does not require the supervision of candidate sets.
@@
In most configurations, RAISE outperforms the compared models in the tasks of generating bottom-right and arbitrary-position answers. To demonstrate the interpretability of RAISE, we interpolate and visualize the learned latent concepts and apply RAISE to solving the odd-one-out task. The experimental results display that RAISE can find the image that breaks the rule on an RPM according to the predictions of concepts in an interpretable way. Finally, we evaluate RAISE on two out-of-distribution configurations, and the results illustrate that RAISE still retains relatively higher accuracy when encountering unseen combinations of rules and attributes.
@@ 2 Related Work
Generative RPM Solvers. We categorize existing RPM solvers as selective and generative solvers in terms of the approach to producing answers. Selective solvers usually choose answers by filling each candidate in the missing position and then scoring the full matrix Zhuo & Kankanhalli (2021); Barrett et al. (2018); Wu et al. (2020); Hu et al. (2021); Benny et al. (2021); Steenbrugge et al. (2018); Hahne et al. (2019); Zhang et al. (2019b); Zheng et al. (2019); Wang et al. (2019; 2020); Jahrens & Martinetz (2020). Generative solvers aim at answer-generation problems by predicting representations or reconstructing images for missing positions Pekar et al. (2020); Zhang et al. (2021b; a). Niv et al. Pekar et al. (2020) extract image representations through Variational AutoEncoder (VAE) Kingma & Welling (2013) and design the relation-wise perception process to infer representations of missing images. By extracting interpretable features of scenes in the perceptron frontend, ALANS Zhang et al. (2021b) and PrAE Zhang et al. (2021a) adopt an algebraic abstract and symbolic logical system as the reasoning backend for target prediction, respectively. These generative solvers realize inspiring generative abstract reasoning abilities to predict answers at the bottom right. LGPP Shi et al. (2021) and CLAP Shi et al. (2023) learn hierarchical latent variables to capture the underlying rules of RPMs with random functions Williams & Rasmussen (2006); Garnelo et al. (2018). They can generate answers at arbitrary positions on RPMs with continuous attributes. RAISE is a variant of conditional latent variable models that achieves generative abstract reasoning on more complex RPMs through rule abstraction and selection in the latent space.
@@
Bayesian Inference with Global Latent Variables. Deep latent variable models (DLVMs) can capture underlying structures in high-dimensional data Kingma & Welling (2013); Sohn et al. (2015); Sønderby et al. (2016). DLVMs can regard shared concepts as global latent variables (e.g., the typical features of elements can act as latent concepts to describe an image set) and introduce local latent variables conditioned on the shared concepts to generate samples with the concepts. GQN Eslami et al. (2018) infers global latent variables to model entire 3D scenes to generate 2D images of unseen perspectives. With object-centric representations Yuan et al. (2023), global latent variables can explain layouts of scenes Jiang & Ahn (2020) or object appearances for multiview scene generation Chen et al. (2021); Kabra et al. (2021); Yuan et al. (2022); Gao & Li (2023); Yuan et al. (in press). Denoting global concepts as the common features of elements, we can represent data with exchange invariance, e.g., sets Edwards & Storkey (2017); Hewitt et al. (2018); Giannone & Winther (2021). NP and its variants Garnelo et al. (2018); Kim et al. (2019); Foong et al. (2020) simulate different function spaces though global latent variables. DLVMs can parse concept-changing rules from RPMs by regarding the rules as global concepts, and generate answers at arbitrary positions Shi et al. (2021; 2023). In this paper, RAISE holds a similar idea of modeling underlying rules as global concepts. Unlike previous works, RAISE attempts to abstract the atomic rules shared among RPMs.
@@ 3 Method
In this paper, an RPM is denoted as $(\bm{x}_{S},\bm{x}_{T})$ where $\bm{x}_{S}$ and $\bm{x}_{T}$ are mutually exclusive, $T$ indexes the target images to predict ($T$ can index multiple target images), and $S$ indexes the context images given as clues. RAISE is trained by maximizing $\log p(\bm{x}_{T}|\bm{x}_{S})$ in the framework of variational Bayes Kingma & Welling (2013). In the generative and inference processes, RAISE abstracts and selects atomic rules in the latent space, which will be introduced in the following sections.
(a) The graphical model of RAISE
(b) The abstract reasoning process
@@ 3.1 Conditional Generation
The generative process is the foundation of answer generation, including the stage of concept learning, abstract reasoning, and image generation.
Concept Learning. RAISE extracts interpretable image representations for abstract reasoning and image generation in the concept learning stage. Previous studies have emphasized the role of abstract object representations in the abstract reasoning of infants Kahneman et al. (1992); Gordon & Irwin (1996), which is similar to the idea of object-centric representation learners that decompose complex scenes into object representations Locatello et al. (2020). Both views reflect the compositionality of human cognition Lake et al. (2011). We can follow similar ideas to convert relatively complex images into simpler latent concepts, and further decompose the overall rules of RPMs into atomic rules shared on the latent concepts. RAISE realizes attribute-level compositionality by encoding attributes of images into $C$ independent latent concepts Wu et al. (2020); Shi et al. (2021; 2023). In the first step, RAISE encodes a context image $\bm{x}_{s}$ into context latent concepts $\bm{z}_{s}=\{\bm{z}_{s}^{c}\}_{c=1}^{C}$:
@@
The encoder $g_{\theta}^{\text{enc}}$ predicts only the means of context latent concepts, and the standard deviation $\sigma_{z}$ is controlled by a hyperparameter to keep training stability. $g_{\theta}^{\text{enc}}$ handles each context image independently, making it possible to extract latent concepts for any set of input images. The encoder does not consider any relationships between images to focus on concept learning.
Abstract Reasoning. As illustrated in Figure 0(b), RAISE predicts the target latent concepts $\bm{z}_{T}$ from context latent concepts $\bm{z}_{S}$ in the abstract reasoning stage, involving rule abstraction, rule selection, and rule execution. To abstract atomic rules and build the global knowledge set, RAISE adopts $K$ groups of global learnable parameters $\psi=\{\psi_{k}\}_{k=1}^{K}$, each indicating an atomic rule that can be shared among RPMs. In the process of rule selection, we use indicators $\{r^{c}\}_{c=1}^{C}$ ($r^{c}=1,...,K$) to select proper rules for each concept out of the global knowledge set $\psi$. The critical of rule selection is to infer the indicators based on $\bm{z}_{S}$. To this end, we first create a 3$\times$3 representation matrix $\bm{Z}^{c}$ for each concept through random target initialization. In the matrix, the representations of context images are filled with the corresponding context latent concepts while those of target images are initialized as zero vectors. Then RAISE extracts the representations for rows and columns:
@@
RAISE averages the row and column representations to obtain $\bm{\bar{p}}^{c}=(\bm{p}_{1}^{c}+\bm{p}_{2}^{c}+\bm{p}_{3}^{c})/3$ and $\bm{\bar{q}}^{c}=(\bm{q}_{1}^{c}+\bm{q}_{2}^{c}+\bm{q}_{3}^{c})/3$, which are respectively overall representations along the row and column. We concatenate $\bm{\bar{p}}^{c}$ and $\bm{\bar{q}}^{c}$ to acquire the probability of selecting rules out of the global knowledge set:
@@
For each concept, the parameter of $h$ is selected from $\psi$ by the indicator $r^{c}$. The neural network $h$ converts the initialized target representations in $\bm{Z}^{c}$ to the means of target latent concepts. Like the concept learning stage, the standard deviation of target latent concepts is controlled by $\sigma_{z}$. RAISE realizes $h$ with convolution layers to predict target latent concepts from the context via message passing on the matrix. Each group of parameters in $\psi$ is the learnable parameters of $h$ that indicate an atomic rule in the global knowledge set. See Appendix C.1 for the detailed description of $h$.
Image Generation. After obtaining target latent concepts in abstract reasoning, RAISE decodes the latent concepts into target images, which is
@@
RAISE generates each target image independently to make the decoder focus on image reconstruction and controls the noise of target images by setting $\sigma_{x}$ as a hyperparameter.

### v3

@@ Abstract
Endowing machines with abstract reasoning ability has been a long-term research topic in artificial intelligence. Raven’s Progressive Matrix (RPM) is widely used to probe abstract visual reasoning in machine intelligence, where models will analyze the underlying rules and select one image from candidates to complete the image matrix. Participators of RPM tests can show powerful reasoning ability by inferring and combining attribute-changing rules and imagining the missing images at arbitrary positions of a matrix. However, existing solvers can hardly manifest such an ability in realistic RPM tests. In this paper, we propose a deep latent variable model for answer generation problems through Rule AbstractIon and SElection (RAISE). RAISE can encode image attributes into latent concepts and abstract atomic rules that act on the latent concepts. When generating answers, RAISE selects one atomic rule out of the global knowledge set for each latent concept to constitute the underlying rule of an RPM. In the experiments of bottom-right and arbitrary-position answer generation, RAISE outperforms the compared solvers in most configurations of realistic RPM datasets. In the odd-one-out task and two held-out configurations, RAISE can leverage acquired latent concepts and atomic rules to find the rule-breaking image in a matrix and handle problems with unseen combinations of rules and attributes.
@@ 1 Introduction
The abstract reasoning ability is pivotal to abstracting the underlying rules from observations and quickly adapting to novel situations Cattell (1963); Zhuo & Kankanhalli (2021); Małkiński & Mańdziuk (2022a), which is the foundation of cognitive processes Gray & Thompson (2004) like number sense Dehaene (2011), spatial reasoning Byrne & Johnson-Laird (1989), and physical reasoning McCloskey (1983). Intelligent systems may benefit from human-like abstract reasoning when leveraging acquired skills in unseen tasks Barrett et al. (2018), for example, generalizing the law of object collision in the simulation environment to real scenes. Therefore, endowing intelligent systems with abstract reasoning ability is the cornerstone of higher-intelligence systems and a long-lasting research topic of artificial intelligence Chollet (2019); Małkiński & Mańdziuk (2022b).
@@
Raven’s Progressive Matrix (RPM) is a classical test of abstract reasoning ability for human and intelligent systems Małkiński & Mańdziuk (2022a), where participators need to choose one image out of eight candidates to fill in the bottom-right position of a 3$\times$3 image matrix Raven & Court (1998). Previous studies demonstrate that participators can display powerful reasoning ability by directly imagining the missing images Hua & Kunda (2020); Pekar et al. (2020), and answer-generation tasks can more accurately reflect the model’s understanding of underlying rules than answer-selection ones Mitchell (2021). For example, some RPM solvers find shortcuts in discriminative tasks by selecting answers according to the bias of candidate sets instead of the given context.
@@
To solve answer-selection problems, many solvers fill each candidate to the matrix for score estimation and can hardly imagine answers from the given context Barrett et al. (2018); Hu et al. (2021). Some generative solvers have been proposed to solve answer-generation tasks Pekar et al. (2020); Zhang et al. (2021b; a). They generate solutions for bottom-right images and select answers by comparing the solutions and candidates. However, some generative solvers do not parse interpretable attributes and attribute-changing rules from RPMs Pekar et al. (2020), and usually introduce artificial priors in the processes of representation learning or abstract reasoning Zhang et al. (2021b; a). On the other hand, most generative solvers are trained with the aid of candidate sets in training, bringing the potential risk of learning shortcuts Hu et al. (2021); Benny et al. (2021).
@@
Deep latent variable models (DLVMs) Kingma & Welling (2013); Sohn et al. (2015) can capture underlying structures of noisy observations via interpretable latent spaces Edwards & Storkey (2017); Eslami et al. (2018); Garnelo et al. (2018); Kim et al. (2019). Previous work Shi et al. (2021) solves generative RPM problems by regarding attributes and attribute-changing rules as latent concepts, which can generate solutions by executing attribute-specific predictive processes. Through conditional answer-generation processes that consider the underlying structure of RPM panels, the distractors are not necessary to train DLVM-based solvers. Although previous work has achieved answer generation in RPMs with continuous attributes, understanding complex discrete rules and abstracting global rules in realistic datasets is still challenging for DLVMs.
@@
This paper proposes a DLVM for generative RPM problems through Rule AbstractIon and SElection (RAISE)  [^1] Code is available at https://github.com/FudanVI/generative-abstract-reasoning/tree/main/raise. RAISE encodes image attributes (e.g., object size and shape) as independent latent concepts to bridge high-dimensional images and latent representations of rules. The underlying rules of RPMs are decomposed into subrules in terms of latent concepts and abstracted into atomic rules as a set of learnable parameters shared among RPMs. RAISE picks up proper rules for each latent concept and combines them into the integrated rule of an RPM to generate the answer. The conditional generative process of RAISE indicates how to use the global knowledge of atomic rules to imagine (generate) target images (answers) interpretably. RAISE can automatically parse latent concepts without meta information of image attributes to reduce artificial priors in the learning process. RAISE can be trained under semi-supervised settings, requiring only a small amount of rule annotations to outperform the compared models in non-grid configurations. By predicting the target images at arbitrary positions, RAISE does not require distractors of candidate sets in training and supports generating missing images at arbitrary and even multiple positions.
@@
RAISE outperforms the compared solvers when generating bottom-right and arbitrary-position answers in most configurations of datasets. We interpolate and visualize the learned latent concepts and apply RAISE in odd-one-out problems to demonstrate its interpretability. The experimental results show that RAISE can detect the rule-breaking image of a matrix through interpretable latent concepts. Finally, we evaluate RAISE on two out-of-distribution configurations where RAISE retains relatively higher accuracy when encountering unseen combinations of rules and attributes.
@@ 2 Related Work
Generative RPM Solvers. While selective RPM solvers Zhuo & Kankanhalli (2021); Barrett et al. (2018); Wu et al. (2020); Hu et al. (2021); Benny et al. (2021); Steenbrugge et al. (2018); Hahne et al. (2019); Zhang et al. (2019b); Zheng et al. (2019); Wang et al. (2019; 2020); Jahrens & Martinetz (2020) focus on answer-selection problems, generative solvers predict representations or images at missing positions Pekar et al. (2020); Zhang et al. (2021b; a). Niv et al. extract image representations through Variational AutoEncoder (VAE) Kingma & Welling (2013) and design a relation-wise perception process for answer prediction Pekar et al. (2020). With interpretable scene representations, ALANS Zhang et al. (2021b) and PrAE Zhang et al. (2021a) adopt algebraic abstract and symbolic logical systems as the reasoning backends. These generative solvers predict answers at the bottom-right position. LGPP Shi et al. (2021) and CLAP Shi et al. (2023) learn hierarchical latent variables to capture the underlying rules of RPMs with random functions Williams & Rasmussen (2006); Garnelo et al. (2018), and can generate answers at arbitrary positions on RPMs with continuous attributes. RAISE is a variant of DLVM to realize generative abstract reasoning on realistic RPM datasets with discrete attributes and rules through atomic rule abstraction and selection.
@@
Bayesian Inference with Global Latent Variables. DLVMs Kingma & Welling (2013); Sohn et al. (2015); Sønderby et al. (2016) can capture underlying structures of high-dimensional data in latent spaces, regard shared concepts as global latent variables, and introduce local latent variables conditioned on the shared concepts to distinguish each sample. GQN Eslami et al. (2018) captures entire 3D scenes via global latent variables to generate 2D images of unseen perspectives. With object-centric representations Yuan et al. (2023), global latent variables can explain layouts of scenes Jiang & Ahn (2020) or object appearances for multiview scene generation Chen et al. (2021); Kabra et al. (2021); Yuan et al. (2022); Gao & Li (2023); Yuan et al. (2024). Global concepts can describe common features of elements in data with exchange invariance like sets Edwards & Storkey (2017); Hewitt et al. (2018); Giannone & Winther (2021). NP family Garnelo et al. (2018); Kim et al. (2019); Foong et al. (2020) constructs different function spaces through global latent variables. DLVMs can generate answers at arbitrary positions of an RPM by regarding the concept-changing rules as global concepts Shi et al. (2021; 2023). RAISE holds a similar idea of modeling underlying rules as global concepts. Unlike previous works, RAISE attempts to abstract the atomic rules shared among RPMs.
@@ 3 Method
In this paper, an RPM problem is $(\bm{x}_{S},\bm{x}_{T})$ where $\bm{x}_{S}$ and $\bm{x}_{T}$ are mutually exclusive sets of images, $S$ indexes the given context images, and $T$ indexes the target images to predict ($T$ can index multiple images). The objective of RAISE is to maximize the log-likelihood $\log p(\bm{x}_{T}|\bm{x}_{S})$ while learning atomic rules shared among RPMs. In the following sections, we will introduce the generative and inference processes of RAISE that can abstract and select atomic rules in the latent space.
(a) Graphical model of RAISE
(b) Abstract reasoning process
@@ 3.1 Conditional Generation
The generative process is the foundation of answer generation, including the stages of concept learning, abstract reasoning, and image generation.
Concept Learning. RAISE extracts interpretable image representations for abstract reasoning and image generation in the concept learning stage. Previous studies have emphasized the role of abstract object representations in the abstract reasoning of infants Kahneman et al. (1992); Gordon & Irwin (1996) and the benefit of disentangled representations for RPM solvers Van Steenkiste et al. (2019), which reflect the compositionality of human cognition Lake et al. (2011). RAISE realizes compositionality by learning latent representations of attributes Shi et al. (2021; 2023). RAISE regards image attributes as latent concepts and decomposes the rules of RPMs into atomic rules based on the latent concepts. Since the description of attributes is not provided in training, the latent concepts learned by RAISE are not exactly the same as the realistic attributes defined in the dataset. RAISE extracts $C$ context latent concepts $\bm{z}_{s}=\{\bm{z}_{s}^{c}\}_{c=1}^{C}$ for each context image $\bm{x}_{s}$ ($s\in S$):
@@
The encoder $g_{\theta}^{\text{enc}}$ outputs the mean of context latent concepts. The standard deviation is controlled by a hyperparameter $\sigma_{z}$ to keep training stability. Each context image is processed through $g_{\theta}^{\text{enc}}$ independently, making it possible to extract latent concepts for any set of input images. In this stage, the encoder does not consider any relationships between images and focuses on concept learning.
Abstract Reasoning. As illustrated in Figure 1(b), RAISE predicts target latent concepts $\bm{z}_{T}$ from context latent concepts $\bm{z}_{S}$ in the abstract reasoning stage, involving rule abstraction, rule selection, and rule execution processes. To abstract atomic rules and build the global knowledge set, RAISE adopts $K$ global learnable parameters $\psi=\{\psi_{k}\}_{k=1}^{K}$, each indicating an atomic rule shared among RPMs. In rule selection, we use categorical indicators $\{r^{c}\}_{c=1}^{C}$ ($r^{c}=1,...,K$) to select a proper rule out of $\psi$ for each concept. Inferring the indicators from $\bm{z}_{S}$ correctly is critical to rule selection. RAISE creates a 3$\times$3 representation matrix $\bm{Z}^{c}$ for each concept, initializing the representations of context images with the corresponding context latent concepts and those of target images with zero vectors. Then RAISE extracts the row-wise and column-wise representations:
@@
RAISE averages the representations via $\bm{\bar{p}}^{c}=(\bm{p}_{1}^{c}+\bm{p}_{2}^{c}+\bm{p}_{3}^{c})/3$ and $\bm{\bar{q}}^{c}=(\bm{q}_{1}^{c}+\bm{q}_{2}^{c}+\bm{q}_{3}^{c})/3$ to obtain integrated representations of row and column rules. We concatenate $\bm{\bar{p}}^{c}$ and $\bm{\bar{q}}^{c}$ to acquire the probability of selecting atomic rules out of the global knowledge set:
@@
RAISE instantiates $h$ by selecting the $r^{c}$-th learnable parameters from the global knowledge set $\psi$ to convert the zero-initialized target representations in $\bm{Z}^{c}$ into the mean of target latent concepts. As in the concept learning stage, the standard deviation of target latent concepts is controlled by $\sigma_{z}$. $h$ consists of convolution layers to aggregate information from neighbor context latent concepts on the matrix and update target latent concepts. Each learnable parameters in $\psi$ indicates a type of atomic rule. See Appendix C.1 for the detailed description of $h$.
Image Generation. Finally, RAISE decodes the target latent concepts predicted in the abstract reasoning stage into the mean of target images:
@@
RAISE generates each target image independently to make the decoder focus on image reconstruction. We control the noise of target images by setting the standard deviation $\sigma_{x}$ as a hyperparameter.